Calculate safe dosages for infants, children, and adults.
Answer #1

The dose is calculated based on the body weight of the patient  .In case of chemotherapy drugs the exact dose is calculated based on the body surface area. The simple step in calculation of safe dose is by

  • measuring the weight of the body in kg from pounds conversion
  • convert the dose in mg
  • Now divide the dose by number of times the drug to be given in a day
  • Now convert the mg to ml
  • For example :A patient weighing 30 pounds has to be given a drug of 1000mg per day in two divide doses. The medication is available as 100mg/ml.Now to calculate the safe dose
    • Convert pound to kg (30×0.454)=13.6 or 14 kg on rounding off
    • Order to give is 500mg
    • Now divide 500mg with the number of tines the drug to be given 500÷2=250mg
    • Available 100mg/ml ,therefore 250mg=?? (250÷100) =2.5ml
    • So to the safe dose is 2.5
  • For example the same drugs has to be given 2mg/kg/day in two divided doses
    • As 2mg=1 kg then 14kg=??
    • 28mg/day (14×2÷1)
    • so one dose is 14mg (28mg/2)
  • This is applicable and most accurate calculations for safe dosages for infant ,child and adult